# Automatic Stage Lighting Control: Is it a Rule-Driven Process or Generative Task?

### v1

@@ 3.4.1. Pre-training
For pre-training the model on music data, we opt for the MLM task, following the approaches in (Zhao et al., 2024b; Zhu et al., 2021; Lewis, 2019). A MLP is used after the decoder’s output to map the output hidden state to music sequence. We first utilize Open L3 to embed the audio music, resulting in an embedding sequence $e=\{e_{1},e_{2},\dots,e_{n}\}\in\mathbf{R}^{n\times h}$, where $h$ is the hidden dimension. We then randomly replace $k\%$ of the tokens with [MASK] tokens and train the model to recover them. Since the music tokens are continuous, we design the [MASK] token as a random variable sampled from a normal distribution, based on (Zhao et al., 2024b):
(4) [MASK] $=[M_{1},M_{2},\dots,M_{h}]^{\prime}\in\mathbb{R}^{1\times h}\ ,$
@@
where represents transposition, $\mu^{i}$ and $\sigma^{i}$ denote the mean and standard deviation of sequence $e$ in the $i^{th}$ feature dimension. In this manner, the model not only aims to recover the obscured tokens but also to identify which tokens have been replaced by [MASK], further assisting the model in learning the structure of the music sequence. To enhance the realism of the recovered output, we also employ a GAN-based discriminator to distinguish between the original music and the reconstructed music, as described by (Zhao et al., 2024b):
[…]
where $D$ and $R$ represent the discriminator and recoverer, respectively. Here, to enhance parameter efficiency, we allow the discriminator and recoverer to use a shared Skip-BART backbone while employing separate heads, as suggested by (Zhao et al., 2024b). Thus, we can express the pre-training loss function of Skip-BART as follows:
(6) $L_{pre}$ $=\alpha_{1}l_{1}+\alpha_{2}l_{2}+\alpha_{3}l_{3}\ ,$
@@
(7) $L_{dis}=\text{CrossEntropy}(D(\hat{e}),0)+\text{CrossEntropy}(D(e),1)\ .$
3.4.2. Fine-tuning
During fine-tuning, we utilize the Language Model (LM) task (Radford et al., 2018) (next token prediction) to supervise the model in generating the corresponding lighting sequences. We frame this task as a classification problem where the model needs to classify the hue and value at each token position. The classification vocabularies are defined by the binning scheme used in the histograms $D_{h,t}$ and $D_{v,t}$, as described in Eq. 1. Consequently, we employ two MLPs to map the decoder output to the classification probabilities for hue and value, respectively. The loss function can be expressed as:
(8) $L_{stf}=\beta_{1}\text{CrossEntropy}(\hat{h},\overline{h})+\beta_{2}\text{CrossEntropy}(\hat{v},\overline{v})\ .$
where $\hat{h}$ and $\hat{v}$ denote the model outputs, while $\beta_{1}$ and $\beta_{2}$ are hyper-parameters. We observe that the learning speeds for these two attributes are not the same. To encourage the model to focus more on the attributes that are learned more slowly, we use adaptive weights as proposed in (Zhao, 2025):
(9) $\beta_{i}^{j}$ $=\frac{\frac{1}{acc_{i}^{j-1}}}{\sum_{i=1}^{2}\frac{1}{acc_{i}^{j-1}}}\ ,$
@@ 3.4.2. Fine-tuning
3.4.3. Inference
During the inference phase, Skip-BART generates the lighting sequence in an auto-regressive manner. To prevent model degradation and to implement a stochastic temperature-controlled sampling method (Hsiao et al., 2021), we represent this as follows:
(10) $\hat{y}_{i}\sim\text{softmax}\left(\frac{\textbf{Skip-BART}(x,[\hat{y}_{1},\dots,\hat{y}_{i-1}])}{t}\right)\ ,$
where $t$ represents the temperature, and $[\hat{y}_{1},\dots,\hat{y}_{i-1}]$ denotes the previously generated lighting sequence. The right term of Eq. 10 consists of two vectors, representing the probabilities of each category for hue and value, respectively. The hue and value are then sampled from these two probability vectors.
@@ 3.4.3. Inference
where $\mathcal{D}(\cdot,\cdot)$ is the distance function $d^{t}$ is the threshold. To satisfy this restriction, we set the probability of all categories over this distance to zero during stochastic temperature-controlled sampling.
4. Evaluation
4.1. Setup
@@
To illustrate the effectiveness of our method, we compare it with other approaches through both quantitative analysis and an human evaluation. In the quantitative analysis, we evaluate the differences among our method, a conventional rule-based method, and several ablation studies in relation to the ground truth. We assume that the closer the generated results are to the ground truth, the better the model performs, a hypothesis that will be validated through human studies. During the human evaluation, participants are asked to score various methods and the ground truth across different dimensions. The experimental subjects can be summarized as follows:
@@
  • Ablation Study: In comparison to conventional BART, our method’s variants primarily focus on the skip connection module, the pre-training and transfer learning mechanisms, and the embedding layer. In the ablation study, we assess the model’s performance without these mechanisms. We denote them as: (i) Skip-BART (w/o skip connection): The model without the skip connection; (ii) Skip-BART (from scratch): The model trained from scratch, without any prior knowledge from transfer learning or pre-training; (iii) Skip-BART (w/o light embedding): The model that does not utilize the proposed discrete light embedding but instead employs a multi-layer perceptron (MLP)-based continuous embedding layer for light.

### v2

@@ 3.3.1 MLM-based Pre-training
For pre-training the model on music data, we opt for the MLM task, following the approaches in (zhao2024mining; zhu2021musicbert; lewis2019bart). A MLP is used after the decoder’s output to map the output hidden state to music sequence. We first utilize Open L3 to embed the audio music, resulting in an embedding sequence $e=\{e_{1},e_{2},\dots,e_{n}\}\in\mathbf{R}^{n\times h}$, where $h$ is the hidden dimension. We then randomly replace $k\%$ of the tokens with [MASK] tokens and train the model to recover them. Since the music tokens are continuous, we design the [MASK] token as a random variable sampled from a normal distribution, based on (zhao2024mining):
[MASK] $=[M_{1},M_{2},\dots,M_{h}]^{\prime}\in\mathbb{R}^{1\times h}\ ,$ (2)
@@
where represents transposition, $\mu^{i}$ and $\sigma^{i}$ denote the mean and standard deviation of sequence $e$ in the $i^{th}$ feature dimension. In this manner, the model not only aims to recover the obscured tokens but also to identify which tokens have been replaced by [MASK], further assisting the model in learning the structure of the music sequence. To enhance the realism of the recovered output, we also employ a GAN-based discriminator to distinguish between the original music and the reconstructed music, as described by (zhao2024mining):
$\underset{R}{\text{min}}\ \underset{D}{\text{max}}\ V(D,R)$ (3)
$=\underset{R}{\text{min}}\ \underset{D}{\text{max}}\ E_{e}[log(D(e))]+E_{e}[log(1-D(R(e)))]\ ,$
where $D$ and $R$ represent the discriminator and recoverer, respectively. For the discriminator, we adopt the same architecture as the sequence classifier in chou2021midibert. It employs a modified self-attention layer to capture the global information from the embedding sequence $e$, followed by a linear layer that classifies the extracted feature as True or False.
Here, to enhance parameter efficiency, we allow the discriminator and recoverer to use a shared Skip-BART backbone while employing separate heads, as suggested by (zhao2024mining). Thus, we can express the pre-training loss function of Skip-BART as follows:
$L_{pre}$ $=\alpha_{1}l_{1}+\alpha_{2}l_{2}+\alpha_{3}l_{3}\ ,$ (4)
@@
$L_{dis}=\text{CrossEntropy}(D(\hat{e}),0)+\text{CrossEntropy}(D(e),1)\ .$ (5)
Due to page limitations, we have included additional details regarding pre-training in Appendix A.
3.3.2 End-to-End Fine-tuning
During fine-tuning, we utilize the Language Model (LM) task (radford2018improving) (next token prediction) to supervise the model in generating the corresponding lighting sequences. We frame this task as a classification problem where the model needs to classify the hue and value at each token position. The classification vocabularies are defined by the binning scheme used in the histograms $D_{h,t}$ and $D_{v,t}$, as described in Eq. 11. Consequently, we employ two MLPs to map the decoder output to the classification probabilities for hue and value, respectively. The loss function can be expressed as:
$L_{stf}=\beta_{1}\text{CrossEntropy}(\hat{h},\overline{h})+\beta_{2}\text{CrossEntropy}(\hat{v},\overline{v})\ .$ (6)
where $\hat{h}$ and $\hat{v}$ denote the model outputs, while $\beta_{1}$ and $\beta_{2}$ are hyper-parameters. We observe that the learning speeds for these two attributes are not the same. To encourage the model to focus more on the attributes that are learned more slowly, we use adaptive weights as proposed in (zhao2024adversarial):
$\beta_{i}^{j}$ $=\frac{\frac{1}{acc_{i}^{j-1}}}{\sum_{i=1}^{2}\frac{1}{acc_{i}^{j-1}}}\ ,$ (7)
@@ 3.3.2 End-to-End Fine-tuning
3.3.3 RSTC-based Inference
During the inference phase, Skip-BART generates the lighting sequence in an auto-regressive manner. To prevent model degradation and to implement a stochastic temperature-controlled sampling method (hsiao2021compound), we represent this as follows:
$\hat{y}_{i}\sim\text{softmax}\left(\frac{\textbf{Skip-BART}(x,[\hat{y}_{1},\dots,\hat{y}_{i-1}])}{t}\right)\ ,$ (8)
where $t$ represents the temperature, and $\hat{y}_{1},\dots,\hat{y}_{i-1}]$ denotes the previously generated lighting sequence. The right term of Eq. 8 consists of two vectors, representing the probabilities of each category for hue and value, respectively. The hue and value are then sampled from these two probability vectors.
@@ 3.3.3 RSTC-based Inference
where $\mathcal{D}(\cdot,\cdot)$ is the distance function $d^{t}$ is the threshold. To satisfy this restriction, we set the probability of all categories over this distance to zero during stochastic temperature-controlled sampling. We refer to this method as Restricted Stochastic Temperature-Controlled (RSTC).
4 Evaluation
4.1 Setup
@@
To illustrate the effectiveness of our method, we compare it with other approaches through both quantitative analysis and an human evaluation. In the quantitative analysis, we evaluate the differences among our method, a conventional rule-based method, and several ablation studies in relation to the ground truth. We assume that the closer the generated results are to the ground truth, the better the model performs, a hypothesis that will be validated through human studies. During the human evaluation, participants are asked to score various methods, including the proposed Skip-BART, the Ablation Study, the previous methods (referred to as the Rule-based Method), and the ground truth across different dimensions.
More details about the model configuration and training process can be found in the Appendix B. The experiments were conducted using the PyTorch framework (paszke1912pytorch) on a server running Ubuntu 22.04.5 LTS, equipped with an Intel(R) Xeon(R) Gold 6133 CPU @ 2.50 GHz, two NVIDIA 4090 GPUs, and one NVIDIA A100 GPU.
Moreover, as mentioned in introduction, there is currently no dataset available for the lighting generation task, which is one reason previous methods have relied on rules. We address this gap by proposing the first stage lighting dataset named ‘RPMC-L2’ (Rock, Punk, Metal, and Core - Livehouse Lighting). This dataset is highly diverse, collected by the authors over the past five years, including contributions from numerous bands, livehouses, and four different capturing devices, covering a wide range of styles such as metalcore, alternative rock, and post-punk. We construct the videos using the proposed method in Appendix C.3 from raw footage. Our dataset includes 699 samples, with lengths ranging from 20 seconds to around 5 minutes, making it a relatively medium-sized dataset in MIR. More details can be found in the Appendix C.
4.2 Quantitative Analysis
Table 1: Quantitative Results: The best result is indicated in bold, and the second best is indicated by underline. This notation will remain consistent in the following tables.
[…]
Hue Value Hue Value Hue Value
Rule-based 48.67 93.39 43.43 86.55 0.50 0.58
Skip-BART 36.13 60.74 28.72 51.27 0.88 2.94
w/o skip connection 36.89 68.33 29.44 58.34 1.15 0.30
w/o light embedding 51.04 67.25 41.50 54.87 0.80 0.70
train from scratch 36.63 67.49 28.83 57.22 0.69 0.53
pre-train w/o random [MASK] 49.97 64.45 42.07 52.63 0.54 1.11
pre-train w/o discriminator 50.40 68.09 41.52 56.54 0.46 1.13
In the quantitative experiment, we compare the generation results of different methods against the ground truth. Since our method utilizes random sampling during training, comparing loss and accuracy has limited significance, similar to most generative works (liang2024pianobart). Therefore, we directly compute four metrics between the generated results and the ground truth: (1) Root Mean Square Error (RMSE) and (2) Mean Absolute Error (MAE), both measuring absolute prediction errors. For the Hue channel, RMSE/MAE are computed on the circular angle differences, ensuring that values like $0^{\circ}$ and $179^{\circ}$ are treated as close rather than far apart, as shown in Eq. 9. (3) $\mathrm{corr}(|\Delta|)$ measures the consistency of change magnitudes by computing the Pearson correlation of the absolute first-order differences between predictions and ground truth. The final value is multiplied by 100. The results are presented in Table 1, where it is evident that our method outperforms all ablation studies and the rule-based method in both hue and value, achieving the lowest error. Since the rule-based method does not directly learn from the ground truth, it may introduce some unfairness. We provide further explanations in the human evaluation section.
4.3 Human Evaluation
To better assess how our results align with human preferences, we designed a questionnaire to collect participants’ feedback on the stage lighting effects created by each method. We recruited 38 respondents through social media platforms and live music venues, with the human evaluation spanning half a month. The participants included 20 males and 18 females aged 18–55 (predominantly 18–30), among whom 9 had professional experience in lighting design, music production, or stage art, and over half were frequent live music attendees. All participants reported normal hearing, normal or corrected-to-normal vision, and no history of color blindness or color weakness. All participants are required to evaluate three music pieces alongside the lights across six metrics (erdmann2025development), with each metric scored from 1 to 7 (the higher the score, the better the evaluation). The lighting conditions were derived from four sources: Ground Truth, Skip-BART, Rule-based Method, and Skip-BART (w/o skip connection). For the ablation study, we specifically evaluate Skip-BART without the skip connection. We considered two main factors: first, the efficiency of these components has been widely demonstrated in other fields; second, increasing the number of comparative methods complicates data collection and may discourage user participation, as longer questionnaires can lead to fatigue and confusion while viewing the same lighting video for a particular piece of music. For simplicity, we will refer to Skip-BART (w/o skip connection) as the Ablation Study in this section and the following tables. Due to page limitations, we outline the primary experimental process and results below, while the Appendix E contains the complete survey questionnaire and the analytical results.
As shown in Table 2 and Table 3, Ground Truth receives the highest overall score (M = 4.5, SD = 0.9), closely followed by our Skip-BART(M=4.3, SD=0.9) and the Skip-BART (w/o skip connection)(M=4.1, SD=0.8), both of which are near the GT level. All three methods significantly outperform the rule-based approach(M=2.7, SD=1.3,p<<0.05), a trend that holds for all the metrics. The only metric where Ground Truth does not achieve the best score is in emotion matching, supporting the perspective that the emotional aspects of music may have a limited connection to lighting control (mcdonald2022illuminating). We also observe that for the rule-based method, the score of the emotion metrics(M=3.1, SD=1.5) is higher than that of other metrics, which validates our reproduction efforts.
Table 2: Human Evaluation Scores: The six metric scores and overall evaluations of the four objects are expressed as Mean ± Standard Deviation (M ± SD) . The bold text represents the best result, while the underlined text indicates the second-best result.
Method Emotion Impact Rhythm Smoothness Atmosphere Surprise Overall In-domain Evaluation Ground Truth 4.50±0.93 4.48±0.99 4.61±0.99 4.62±1.07 4.49±0.89 4.34±1.10 4.51±0.88 Skip-BART 4.69±0.87 4.39±0.95 4.50±1.06 4.32±1.12 4.32±0.93 3.83±1.06 4.35±0.87 Ablation Study 4.31±0.94 3.78±0.96 4.54±1.08 4.43±1.12 4.11±0.98 3.50±1.00 4.11±0.84 Rule-based 3.12±1.52 2.65±1.39 2.54±1.47 2.56±1.27 2.77±1.50 2.35±1.40 2.67±1.29 Cross-domain Evaluation Skip-BART 4.31±1.05 4.18±1.09 4.60±1.14 4.61±0.92 4.41±1.09 3.91±1.17 4.34±0.92 Ablation Study 3.94±1.31 3.93±1.12 4.44±1.02 4.42±1.05 4.04±1.11 3.79±1.15 4.10±0.97 Rule-based 3.61±1.31 3.04±1.20 3.21±1.31 3.88±1.15 3.34±1.19 2.96±1.41 3.34±1.13
[…]
Comparison In-domain Evaluation Cross-domain Evaluation
[…]
Ground Truth vs. Skip-BART 0.16 0.10 0.724
Ground Truth vs. Ablation Study 0.40 0.10 0.003**
Ground Truth vs. Rule-based 1.84 0.21 <<0.001**
Skip-BART vs. Ablation Study 0.23 0.10 0.152 0.24 0.12 0.167
Skip-BART vs. Rule-based 1.68 0.19 <<0.001** 1.00 0.17 <<0.001**
Ablation Study vs. Rule-based 1.44 0.16 <<0.001** 0.76 0.15 <<0.001**
Additionally, to evaluate our model’s performance in a cross-domain scenario, we compare different methods across various music styles. For fairness, we utilize songs generated by Suno (suno) instead of selecting them manually. This approach allows us to assess the model’s performance on entirely novel songs. Furthermore, since no ground truth is available, we omit this object and ask 30 users to evaluate other three objects across the following music genres: folk, R&B, and jazz. The detailed results are presented in Tables 2 and 3. Our proposed method consistently achieves the highest scores across all metrics, demonstrating significant differences compared to the rule-based method.
5 Limitations and Future Work
While our framework demonstrates positive results, there remain several limitations and opportunities for further improvement. Below we highlight several key directions worth further exploration:
  • Generalizability Across Music Genres: In our experiment, we have demonstrated promising results from our method on both in-domain music genres (e.g., rock, punk, metal, and hardcore) and cross-domain genres (e.g., folk, R&B, and jazz). However, further evaluation across a wider range of styles could be explored in future research. In practice, our method is independent of the dataset, allowing users to retrain or fine-tune our model based on their specific datasets. However, the scarcity of large, high-quality datasets has long posed a challenge in the MIR community. In addition to the pre-training and transfer learning mechanisms used in this paper, unsupervised domain alignment methods also present a potential paradigm, requiring only the music in the target domain without the need for light ground truth.
  • Alternative Generation Technology Exploration: This paper introduces a generative model for lighting control, framing it as a creative task. However, as discussed in Appendix D, the model occasionally shows overly strong local fluctuations, which suggest that other popular generative paradigms and technologies deserve exploration. Additionally, emerging techniques such as Reinforcement Learning from Human Feedback (RLHF) (ouyang2022training) and transferring from general music foundation models (suno; yuan2025yue) may provide promising directions for enhancing model capability.
  • Towards More Practical Application Scenarios: Similar to most previous studies, our method currently supports only offline primary lighting generation. Future research could explore real-time and multi-light control. However, these upgrades face significant challenges. For multi-stage control, the heterogeneity among venues must be considered, as different livehouses often have varying light positions and quantities. Regarding online control, the primary challenge is the need to predict changes in music quickly and accurately, as light control decisions must be made within seconds. However, accurately predicting a song’s progression is difficult due to its artistic nature, where unexpected elements contribute to the charm of music.
  • Richer Modality Incorporation: In our model, it is also possible to incorporate richer modalities, such as rhythm, musical sections (like verses and choruses), and lyrics. Models such as beat tracking (chang2024beast) and melody extraction (section classification) (zhao2024adversarial) can also be integrated to provide more informative signals.
6 Conclusion
In this paper, we propose the first end-to-end deep learning method, Skip-BART, for stage lighting generation. We modify the BART model to enable the encoder to extract music features in a single pass and the decoder to generate light hue and intensity auto-regressively, framing ASLC as a generative task. A skip connection mechanism is introduced to enhance the relationship between corresponding music and light frames. Additionally, we incorporate effective deep learning techniques to allow the model to learn from limited data, including transfer learning, pre-training, and RSTC sampling. To support research in this field, we present a dataset construction method based on raw video and release the first stage lighting dataset, RPMC-L2. Through quantitative analysis and human studies, we demonstrate the efficiency of our method and highlight the limitations of conventional rule-based approaches. Our Skip-BART shows performance closely matching that of real human lighting engineers. The results validate our perspective that music lighting control resembles an art generation task rather than a simple classification and mapping task based on predefined rules. This study aims to provide deeper insights and support for future research in this area.
Ethics Statement
Our study did not use any privacy data containing personally identifiable information; all information from human evaluation participants was anonymized, and only essential background data were collected for statistical and analytical purposes. At the same time, we recognize that the generated lighting may involve high flicker frequencies, strong color contrasts, or rapid brightness changes, which could pose health or discomfort risks to individuals with photosensitive epilepsy or visual sensitivities.
Reproducibility Statement
The source code, trained parameters, and processed dataset are available in the supplementary materials. Moreover, we outline our data processing methods, training details and evaluation methods in Appendix.
Appendix
Appendix A Pre-training Details
In this section, we aim to provide additional details regarding the input and output, the neural network architecture, and the hyper-parameters involved in the pre-training phase, ensuring reproducibility.
During pre-training, in order to avoid any potential information leakage, we utilize only the music data without light. To provide the decoder with a robust starting point for fine-tuning, we adopt a BART-based MLM approach. In this manner, the encoder encodes the masked music sequence, while the decoder receives the shift-right ground truth as input and attempts to accurately recover the original music sequence. Although there may be a gap between the music modality in pre-training and the light modality in fine-tuning, this pre-training method has been demonstrated to be effective, as supported by the theory of cross-modal transfer learning (niu2021decade), and it has been successfully applied in numerous domains (chen2024deep).
Specifically, in each epoch, for a given embedding sequence $e=\{e_{1},e_{2},\ldots,e_{n}\}$, we first randomly mask some of the embeddings to generate the input for the encoder, denoted as $\tilde{e}=\{\tilde{e}_{1},\tilde{e}_{2},\ldots,\tilde{e}_{n}\}$. If $\tilde{e}_{i}$ is chosen, it is replaced with a random [MASK] token. Otherwise, it retains its original value $e_{i}$. For the decoder, we use the shift-right ground truth as input, $\{[\text{SOS}],e_{1},e_{2},\ldots,e_{n-1}\}$, which outputs the hidden features $f=\{f_{1},f_{2},\ldots,f_{n}\}$. These features are then fed into a regressor (a simple three-layer MLP with LeakyReLU as the activation function) to generate the estimated input sequence, $\hat{e}=\{\hat{e}_{1},\hat{e}_{2},\ldots,\hat{e}_{n}\}$. Additionally, following the recommendations of zhao2024mining, we introduce a discriminator to enhance the realism of the recovered music sequence. The discriminator takes the original music sequence $e$ and the recovered sequence $\hat{e}$ as input, attempting to differentiate between them, while the decoder works to obfuscate this distinction. For the structure of the discriminator, we utilize a sequence-level classifier from MidiBERT (chou2021midibert), which employs a modified self-attention mechanism:
$M_{\text{attn}}$ $=\text{Softmax}(\text{LinearLayer}(e))\in\mathbb{R}_{+}^{n\times d}\,$, (10)
$z$ $=\text{Flatten}(M_{\text{attn}}^{T}\cdot e)\in\mathbb{R}^{hd}\,$,
$y$ $=\text{Softmax}(\text{LinearLayer}(z))\in\mathbb{R}_{+}^{2}\,$,
where $M_{\text{attn}}$ is the attention matrix, $h$ represents the feature dimension of the input, $d$ is the hidden dimension, and $y$ yields a binary classification result.
Lastly, regarding the hyperparameters, we set the mask ratio as a random number sampled from $U(0.15,0.30)$ in each epoch. The weights $\beta_{1},\beta_{2},\text{ and }\beta_{3}$ in the loss function (Eq. 4) are set to 0.8, 0.2, and 0.1, respectively.
Appendix B Detailed Experiment Setup
The experimental subjects can be summarized as follows:
@@ Appendix B Detailed Experiment Setup
  • Ablation Study: In comparison to conventional BART, our method’s variants primarily focus on the skip connection module, the pre-training and transfer learning mechanisms, and the embedding layer. In the ablation study, we assess the model’s performance without these mechanisms. We denote them as: (i) Skip-BART (w/o skip connection): The model without the skip connection; (ii) Skip-BART (from scratch): The model trained from scratch, without any prior knowledge from transfer learning or pre-training; (iii) Skip-BART (w/o light embedding): The model that does not utilize the proposed discrete light embedding but instead employs a MLP-based continuous embedding layer for light; (iv) Skip-BART (pre-train w/o random [MASK]): During pre-training, the [MASK] token is fixed instead of being randomly assigned as defined in Eq. 2; (v) Skip-BART (pre-train w/o discriminator): The model that does not utilize a discriminator during pre-training, i.e., $\beta_{3}$ in Eq. 4 is set to 0.